# Millimeter emission from supermassive black hole coronae

### v1

@@ 1 Introduction
An Active Galactic Nucleus (AGN) is produced when a supermassive black hole (SMBH) accretes matter from its surrounding medium. The material around a SMBH is pulled by the strong gravitational field and spirals inward, forming an accretion disc. Processes related to turbulent magnetic viscosity in the accretion disc cause the material to lose angular momentum and fall towards the SMBH, releasing huge amounts of gravitational energy. The material then heats up, and radiates across a broad range of wavelengths (e.g. Balbus & Hawley, 1998). The accretion process is very efficient in converting gravitational potential energy into other forms, such as radiation, magnetic fields, and particle acceleration. Moreover, close to the innermost regions of the accretion disc, magnetic reconnection events can lead to the heating of extremely hot gas ($\sim$109 K) in a compact region known as a corona (e.g. Kamraj et al., 2022).
Most ($\sim$90%) AGN emit only weakly at radio wavelengths and are therefore classified as radio-quiet (RQ). Although these AGNs lack a powerful radio jet (Urry & Padovani, 1995), radio emission at 22 GHz is commonly detected (Smith et al., 2020; Magno et al., 2025). Interestingly, millimetre (mm) emission is also frequently observed in nearby RQ AGN, with most of the mm continuum flux originating from an unresolved core. Several studies have argued that this radiation is synchrotron emission produced in the corona, which also produces the hard X-ray emission above 2 keV through inverse Compton scattering of thermal disc photons, ubiquitously observed in AGN (e.g. Laor & Behar, 2008; Inoue & Doi, 2014; Behar et al., 2018; Panessa et al., 2019). Recent studies strongly support this interpretation: high-resolution ($\sim$0.1–1”) ALMA observations of nearby AGN revealed a tight correlation between X-ray emission and mm continuum at $\sim$200 GHz (Kawamuro et al., 2022) and at $\sim$100 GHz (Ricci et al., 2023), as well as a correlation between the SMBH mass and the mm flux density at $\sim$230 GHz (Ruffa et al., 2024). Thus, the corona might play a key role in the mm emission from RQ AGN.
@@
In addition, the presence of ultra-relativistic particles in coronae is supported by both theoretical and observational evidence. Magnetic reconnection events, as well as other processes related to internal shocks and/or turbulence, can be responsible for accelerating relativistic particles out of thermal equilibrium (non-thermal particles; e.g. Inoue et al., 2008; Beloborodov, 2017; Sironi & Beloborodov, 2020; Grošelj et al., 2024; Nättilä, 2024). In the case of a corona, the presence of extremely hot gas ($kT>100$ keV) leads to the co-existence of both thermal and non-thermal relativistic particles (e.g. Özel et al., 2000), which we refer to as a ‘hybrid’ plasma. Another independent proof of a non-thermal electron population in the corona is the hard X-ray/soft gamma-ray tail detected in the XRBs Cyg X-1 (Zdziarski et al., 2021) and MAXI J1820+070 (Cangemi et al., 2021). A scenario involving (hadronic) cosmic rays in the corona is also compatible with the MeV polarised emission detected in Cyg X-1 (Romero et al., 2014), and possibly with its sub-GeV and TeV emission (Fang et al., 2024). In addition, $\gamma$-ray emission in the 1–300 GeV energy range was detected in the stacked signal of 37 RQ AGN, which was interpreted to be produced in extended corona (The Fermi-LAT Collaboration et al., 2025). Furthermore, several studies have suggested that significant neutrino emission can be produced by cosmic rays in SMBH coronae (e.g. Inoue et al., 2019; Murase et al., 2020; Kun et al., 2024; Fang et al., 2024), consistent with the neutrino emission from NGC 1068 (Inoue et al., 2020; Eichmann et al., 2022). This scenario is particularly attractive because it can account for gamma-ray emission orders of magnitude lower than the neutrino emission, as the corona environment produces strong gamma-gamma absorption.
@@ 2 Emission model
The nuclear region of a RQ galaxy (inner $\sim$100 pc) can contain large reservoirs of ionised gas, dust, cosmic rays, and an AGN. Thus, the observed continuum radio-to-submm SED of a RQ galaxy is the combination of several contributions, each with a distinct shape. We assume that these components are independent of each other and that the total emission is simply the sum of all components. The most relevant SED components that we consider are [^1] All frequencies are given in the rest frame. For spectral indices we adopt the convention $S_{\nu}\propto\nu^{\alpha}$.:
@@
    where $A_{\mathrm{d}}$ is a normalisation constant, $\nu_{0}$ is a reference frequency (fixed at 100 GHz), and $\tau_{\mathrm{d}}=(\nu/\nu_{\tau_{1}})^{\beta}$. We note that the dust mass is $M_{\mathrm{d}}\propto\nu_{\tau_{1}}$ (Draine, 2011), so that an accurate determination of $\nu_{\tau_{1}}$ can help to estimate dust masses, although this would typically require a good spectral coverage between 400–1 000 GHz that is challenging to obtain. We also note that, at the frequencies at which the dust becomes optically thick, it could also absorb the emission from the corona. However, this would depend on geometrical assumptions regarding the dust distribution along the line of sight, and it would only affect the optically thin part of the corona SED, which is difficult to observe as it is expected to be faint and typically below the dust continuum emission. We thus neglect this effect in our model fitting. In addition, rapidly spinning dust grains can produce Anomalous Microwave Emission (AME) that peaks around $\nu\sim 30$ GHz, although in certain conditions, it can reach $\sim$100 GHz (Dickinson et al., 2018; Murphy et al., 2020). However, this component is faint and, in extragalactic contexts, it has only been detected in off-nuclear star-forming regions in a handful of nearby galaxies (e.g. Poojon et al., 2024). It is thus unlikely that this component is bright in the sources of our sample, which in fact do not present a bump in their SED at (rest-frame) frequencies of $\sim$30 GHz (Sect. 3). For this reason, we do not include an AME emission component in our model.
@@ 2.1 Corona emission
A related parameter commonly used to characterise the plasma conditions is the $\sigma$ parameter defined as $\sigma=B^{2}/(4\pi n_{\mathrm{th,e}}m_{\mathrm{e}}c^{2})$. Neglecting some factors of order unity, one can approximate $\sigma\approx\epsilon_{B}\,\Theta$, which helps to understand the dependence of $\sigma$ on the model parameters. In particular, for a highly magnetised plasma, one expects $\sigma$ to be of order of unity.
@@ 2.2 Parameter fitting
We developed a code [^3] The code will be made publicly available in https://github.com/santimda/BH-corona_emission that calculates the total SED from the physical coronal model, together with the phenomenological components described in Sect. 2. We use the Python package Bilby (Ashton et al., 2019) for fitting the SED using the standard sampler for Markov Chain Monte Carlo emcee (Foreman-Mackey et al., 2013). One key issue is dealing with observations with different angular resolutions. The corona is extremely compact, so its emission should be point-like in all observations. The other galactic components are most likely extended and inhomogeneous, so it is not possible to scale the observed flux density with the beam size in a reliable way. For this reason, we use beam-matched data whenever possible, even if that requires compromising angular resolution. The main drawback is that the corona emission is not very bright, and so it can be well below the continuum of other galactic components in observations with poor angular resolution that collect too much diffuse emission, making it undetectable. For this reason, we filter out observations with poor angular resolution (typically $>1"$, corresponding to linear scales $>$100 pc for galaxies at distances beyond 20 Mpc) and consider them only as strict upper limits (ULs) for the total flux density. Similarly, when it comes to observations with significantly higher angular resolution than the rest of the dataset (e.g., from very large baseline interferometry—VLBI), we consider them only as strict lower limits (LLs) of the total flux density. In both cases, we leave a 1-$\sigma$ margin to be more conservative about using points as strict ULs or LLs, that is, for a measured flux of $S_{\nu}\pm\Delta S_{\nu}$, we choose $S_{\mathrm{UL}}=S_{\nu}+\Delta S_{\nu}$ and $S_{\mathrm{LL}}=S_{\nu}-\Delta S_{\nu}$.
@@ 3.3 Analysis of the sample sources
We summarize the results of the SED fitting of the two relevant corona parameters, $r_{\mathrm{c}}$ and $\delta$, in Fig. 6. The parameters range in the values $r_{\mathrm{c}}=60$–250 and $\delta=10^{-3}$–10-1, with median values of $r_{\mathrm{c}}\approx 140$ and $\log{\delta}\approx-2$ ($\delta\approx 1$%). The derived values of $B$ range between 10–150 G, with a median of $\approx$15 G. We also find values of $\sigma$ ranging between 0.01–1, with a mean value of $\lesssim$0.1, roughly consistent with the assumption of a significantly magnetised medium ($\sigma\sim 1$). We highlight that the value we obtain of $\delta\approx 1$% is roughly in agreement with the expected value of $\sim$3–4suggested by Inoue et al. (2008) based on the requirements to explain the MeV background.
Figure 6: Non-thermal particle content and corona sizes obtained for the sources in the sample via the SED fitting (with the parameters defined as in Sects. 2.1 and 2.2). The grey dashed lines mark the median values $r_{\mathrm{c}}=140$ and $\log{\delta}=-2$. For the sources GRS 1734--292 and IC 4329A, the broad range in $r_{\mathrm{c}}$ reflects the range in sizes derived in different epochs.
@@
Figure 8: Correlation between the luminosity at 220 GHz and the black hole mass. The points and the blue curve are from Ruffa et al. (2024); filled circles are sources with dynamical $M_{\mathrm{BH}}$ measurements and open squares sources with $M_{\mathrm{BH}}$ derived using the $M_{\mathrm{BH}}$$\sigma_{\star}$ relation (Ruffa et al., 2024). Upper limits are plotted in green for easier distinction. We also plot with pink stars the sources studied in this work (except NGC 3227, which was already included in the sample). Colour-gradient curves are the corona luminosities calculated for the model parameters defined in Sect. 3.1, and using the mean values of $r_{\mathrm{c}}=130$ and $\delta=-2.0$ (Sect. 3.3). We explore different values of $r_{\mathrm{c}}$ (top panel) and $\delta$ (bottom panel), covering the ranges derived for the sources in the sample.
@@ 4 Discussion
We carried out an analysis of the SEDs of seven RQ AGNs using a multi-component SED fitting scheme. Here, we provide an overview of the main limitations of this approach, both from the modelling perspective and the observational aspects, in order to suggest possible ways to move forward.
@@
First, we highlight that we are using a one-zone coronal emission model. Even though the assumption of a homogeneous emitter is most likely an oversimplification, it still yields results that are currently consistent with the observations. It remains extremely difficult to constrain more complex models (e.g. Raginski & Laor, 2016) unless the optically thick part of the radio-mm SED is clearly detected. In most sources, this will require very deep and high-angular-resolution observations at frequencies below 100 GHz. Notably, the correlation between radio and X-ray luminosity is weaker at 22 GHz (Magno et al., 2025) than at 100 GHz (Ricci et al., 2023), consistent with the weaker coronal emission at 22 GHz predicted by our one-zone model. Another limitation of assuming a homogeneous spherical corona is that we cannot predict realistic images of it, although their characteristic sizes of below a few $\mu$as make them unresolvable with current facilities. As a final note on the corona, we note that when observing very dust-obscured sources, we could expect absorption of its emission by dust at frequencies above $\nu_{\tau_{1}}$, although this depends on the dust being along our line of sight to the corona.
@@
Further limitations of the methodology used here include not accounting for an AME component, which could be relevant in sources that present a bump in the SED at $\sim$30 GHz. This component could be added at the expense of additional free parameters. Regarding the synchrotron component, we preferred to ignore the effects of f–f absorption (FFA) in this component due to the limited data at high angular resolution at low frequencies, although this possibility is available in our code (an application of a clumpy absorption medium was presented in Mutie et al., 2025). In this regard, it could be possible to adopt a more model-dependent approach in which, under certain assumptions about the diffuse gas distribution, the FFA is linked to the f–f emission, but this should be properly tested on a large sample of galaxies not necessarily limited to RQ AGN with corona emission. For this reason, we currently favour a more phenomenological approach in which the parameters that describe the FFA and f–f emission are not tied to further assumptions.
@@
Finally, we highlight the observational challenges involved in studying the corona emission of RQ AGN. From a fitting perspective, it is crucial to have a well-sampled SED from radio to FIR with consistent angular resolutions and $uv$-coverage. This ensures that the same scales are probed, although it is very challenging to obtain such a consistent dataset (so far only attempted in Mutie et al., 2025). High-resolution IR imaging is currently unavailable, which hinders our chances of constraining the dust component. The next generation VLA (ngVLA [^8] https://ngvla.nrao.edu) is very promising due to the extremely good angular resolution and sensitivity to frequencies up to $\sim$100 GHz that it will provide. In the mm to sub-mm domain, ALMA will remain the most suitable observatory. One of the most difficult issues to address in the SED fitting is the corona variability on timescales of days, although to date there are only very limited studies of this phenomenon. In this regard, the prospects for the ALMA Wide Sensitivity Upgrade are particularly promising, as it will allow us to obtain a broad bandwidth in a single epoch, something particularly useful around $\sim$100 GHz (as shown in Shablovinskaya et al., 2024). Complementing with simultaneous hard X-ray observations is ideal, as it can further reduce the uncertainty in the corona parameters $\tau_{\mathrm{T}}$ and $k\,T_{\mathrm{c}}$, although this should not be a deciding factor in most cases (Fig. 2).

### v2

@@ 1 Introduction
An Active Galactic Nucleus (AGN) is produced when a supermassive black hole (SMBH) accretes matter from its surrounding medium. The material around an SMBH is pulled by the strong gravitational field and spirals inward, forming an accretion disc. Processes related to turbulent magnetic viscosity in the accretion disc cause the material to lose angular momentum and fall towards the SMBH, releasing huge amounts of gravitational energy. The material then heats up, and radiates across a broad range of wavelengths (e.g. Balbus & Hawley, 1998). The accretion process is very efficient in converting gravitational potential energy into other forms, such as radiation, magnetic fields, and particle acceleration. Moreover, close to the innermost regions of the accretion disc, magnetic reconnection events can heat the gas to extremely high temperatures ($\sim$109 K), giving rise to a compact, hot region known as the corona (e.g. Haardt & Maraschi, 1991; Kamraj et al., 2022).
Most ($\sim$90%) AGN emit only weakly at radio wavelengths and are therefore classified as radio-quiet (RQ). Although these AGNs lack a powerful radio jet (Urry & Padovani, 1995), radio emission at 22 GHz is commonly detected (Smith et al., 2020; Magno et al., 2025). Interestingly, millimetre (mm) emission is also frequently observed in nearby RQ AGN, with most of the mm continuum flux originating from an unresolved core. Several studies have argued that this radiation is synchrotron emission produced in the corona, which also produces the hard X-ray emission above 2 keV through inverse Compton scattering of thermal disc photons, ubiquitously observed in AGN (e.g. Laor & Behar, 2008; Inoue & Doi, 2014; Behar et al., 2018; Panessa et al., 2019). Recent studies strongly support this interpretation: high-resolution ($\sim$0.1–1”) ALMA observations of nearby AGN revealed a tight correlation between X-ray emission and mm continuum at $\sim$200 GHz (Kawamuro et al., 2022) and at $\sim$100 GHz (Ricci et al., 2023), as well as a correlation between the SMBH mass and the mm flux density at $\sim$230 GHz (Ruffa et al., 2024). In addition, correlated variability of X-ray and mm emission on year-timescales was found in NGC 1566 Jana et al. (2025). Together, these results support the idea that the corona plays a key role in the mm emission from RQ AGN.
@@
In addition, the presence of ultra-relativistic particles in coronae is supported by both theoretical and observational evidence. Magnetic reconnection events, as well as other processes related to internal shocks and/or turbulence, can be responsible for accelerating relativistic particles out of thermal equilibrium (non-thermal particles; e.g. Inoue et al., 2008; Beloborodov, 2017; Sironi & Beloborodov, 2020; Grošelj et al., 2024; Nättilä, 2024). In the case of a corona, the presence of extremely hot gas ($kT>100$ keV) leads to the co-existence of both thermal and non-thermal relativistic particles (e.g. Özel et al., 2000), which we refer to as a ‘hybrid’ plasma. Another independent proof of a non-thermal electron population in the corona is the hard X-ray/soft gamma-ray tail detected in the XRBs Cyg X-1 (Zdziarski et al., 2021) and MAXI J1820+070 (Cangemi et al., 2021). A scenario involving (hadronic) cosmic rays in the corona is also compatible with the MeV polarised emission detected in Cyg X-1 (Romero et al., 2014), and possibly with its sub-GeV and TeV emission (Fang et al., 2024). In addition, $\gamma$-ray emission in the 1–300 GeV energy range was detected in the stacked signal of 37 RQ AGN, which was interpreted to be produced in extended coronae (The Fermi-LAT Collaboration et al., 2025). Furthermore, several studies have suggested that significant neutrino emission can be produced by cosmic rays in SMBH coronae (e.g. Inoue et al., 2019; Murase et al., 2020; Kun et al., 2024; Fang et al., 2024), consistent with the neutrino emission from NGC 1068 (Inoue et al., 2020; Eichmann et al., 2022). This scenario is particularly attractive because it can account for gamma-ray luminosity orders of magnitude lower than the neutrino luminosity, as the corona environment produces strong gamma-gamma absorption.
@@ 2 Emission model
The nuclear region of an RQ galaxy (inner $\sim$100 pc) can contain large reservoirs of ionised gas, dust, cosmic rays, and an AGN. Thus, the observed continuum radio-to-submm SED of an RQ galaxy is the combination of several contributions, each with a distinct shape. We assume that these components are independent of each other and that the total emission is simply the sum of all components. The most relevant SED components that we consider are [^1] All frequencies are given in the rest frame. For spectral indices we adopt the convention $S_{\nu}\propto\nu^{\alpha}$.:
@@
    where $A_{\mathrm{d}}$ is a normalisation constant, $\nu_{0}$ is a reference frequency (fixed at 100 GHz), and $\tau_{\mathrm{d}}=(\nu/\nu_{\tau_{1}})^{\beta}$. We note that the dust mass is $M_{\mathrm{d}}\propto\nu_{\tau_{1}}$ (Draine, 2011), so that an accurate determination of $\nu_{\tau_{1}}$ can help to estimate dust masses, although this would typically require a good spectral coverage between 400–1 000 GHz that is challenging to obtain. We also note that, at the frequencies at which the dust becomes optically thick, it could also absorb the emission from the corona. However, this would depend on geometrical assumptions regarding the dust distribution along the line of sight, and it would only affect the optically thin part of the corona SED, which is difficult to observe as it is expected to be faint and typically below the dust continuum emission. We thus neglect this effect in our model fitting. In addition, rapidly spinning dust grains can produce Anomalous Microwave Emission (AME) that peaks around $\nu\sim 30$ GHz, although in certain conditions, it can reach $\sim$100 GHz (Dickinson et al., 2018; Murphy et al., 2020). However, this component is faint and, in extragalactic contexts, it has only been detected in off-nuclear star-forming regions in a handful of nearby galaxies (e.g. Poojon et al., 2024; Fernández-Torreiro et al., 2024). It is thus unlikely that this component is bright in the sources of our sample, which in fact do not present a bump in their SED at (rest-frame) frequencies of $\sim$30 GHz (Sect. 3). For this reason, we do not include an AME emission component in our model.
@@ 2.1 Corona emission
A related parameter commonly used to characterise the plasma conditions is the $\sigma$ parameter defined as $\sigma=B^{2}/(4\pi n_{\mathrm{th,e}}m_{\mathrm{e}}c^{2})$. Neglecting some factors of order unity, one can approximate $\sigma\approx\epsilon_{B}\,\Theta$, which helps to understand the dependence of $\sigma$ on the model parameters. In particular, for a highly magnetised plasma, one expects $\sigma$ to be of the order of unity. Similarly, this can be compared with the plasma $\beta_{B}$ parameter defined as the ratio of thermal and magnetic pressures, $\beta_{B}=P_{\mathrm{th}}/P_{B}$, which can be rewritten as [^3] If a two-temperature plasma is assumed for the corona, for $r_{\mathrm{c}}\sim 140$ we expect an ion temperature of $T_{\mathrm{i}}\sim 1.1$ MeV (e.g. Eq. 1 in Inoue et al., 2024), and thus a total thermal pressure of $P_{\mathrm{th}}\sim(1+T_{\mathrm{i}}/T_{\mathrm{c}})\,U_{\mathrm{th,e}}/a(\Theta)\sim 5U_{\mathrm{th,e}}$. $\beta_{B}\sim 5/\epsilon_{B}$.
@@ 2.2 Parameter fitting
We developed a code [^4] https://github.com/santimda/sed_fitter_BHcorona that calculates the total SED from the physical coronal model, together with the phenomenological components described in Sect. 2. We use the Python package Bilby (Ashton et al., 2019) for fitting the SED using the standard sampler for Markov Chain Monte Carlo emcee (Foreman-Mackey et al., 2013). One key issue is dealing with observations with different angular resolutions. The corona is extremely compact, so its emission should be point-like in all observations. The other galactic components are most likely extended and inhomogeneous, so it is not possible to scale the observed flux density with the beam size in a reliable way. For this reason, we use beam-matched data whenever possible, even if that requires compromising angular resolution. The main drawback is that the corona emission is not very bright, and so it can be well below the continuum of other galactic components in observations with poor angular resolution that collect too much diffuse emission, making it undetectable. For this reason, we filter out observations with coarse angular resolution (typically $>1"$, corresponding to linear scales $>$100 pc for galaxies at distances beyond 20 Mpc) and consider them only as strict upper limits (ULs) for the total flux density. Similarly, when it comes to observations with significantly higher angular resolution than the rest of the dataset (e.g., from very large baseline interferometry—VLBI), we consider them only as strict lower limits (LLs) of the total flux density. In both cases, we leave a 1-$\sigma$ margin to be more conservative about using points as strict ULs or LLs, that is, for a measured flux of $S_{\nu}\pm\Delta S_{\nu}$, we choose $S_{\mathrm{UL}}=S_{\nu}+\Delta S_{\nu}$ and $S_{\mathrm{LL}}=S_{\nu}-\Delta S_{\nu}$.
@@ 3.3 Analysis of the sample sources
We summarize the results of the SED fitting of the two relevant corona parameters, $r_{\mathrm{c}}$ and $\delta$, in Fig. 6. The parameters range in the values $r_{\mathrm{c}}=60$–250 and $\delta=10^{-3}$–10-1, with median values of $r_{\mathrm{c}}\approx 133$ and $\log{\delta}\approx-1.86$ ($\delta\approx 1.4$%). The derived values of $B$ range between 10–150 G, with a median of $\approx$20 G. We also find values of $\sigma$ ranging between 0.01–1, with a mean value of $\lesssim$0.1, roughly consistent with the assumption of a significantly magnetised medium ($\sigma\sim 1$).
Figure 6: Non-thermal particle content and corona sizes obtained for the sources in the sample via the SED fitting (with the parameters defined as in Sects. 2.1 and 2.2). The grey dashed lines mark the median values $r_{\mathrm{c}}=133$ and $\log{\delta}=-1.86$. For the sources GRS 1734--292 and IC 4329A, the broad range in $r_{\mathrm{c}}$ reflects the range in sizes derived in different epochs.
@@
Figure 8: Correlation between the luminosity at 220 GHz and the black hole mass. The points and the blue curve are from Ruffa et al. (2024); filled circles are sources with dynamical $M_{\mathrm{BH}}$ measurements and open squares sources with $M_{\mathrm{BH}}$ derived using the $M_{\mathrm{BH}}$$\sigma_{\star}$ relation (Ruffa et al., 2024). Upper limits are plotted in green for easier distinction. We also plot with pink stars the sources studied in this work (except NGC 3227, which was already included in their sample). Colour-gradient curves are the corona luminosities calculated for the model parameters defined in Sect. 3.1, and using the approximate mean values of $r_{\mathrm{c}}=130$ and $\log{\delta}=-2.0$ (Sect. 3.3). We explore different values of $r_{\mathrm{c}}$ (top panel) and $\delta$ (bottom panel), covering the ranges derived for the sources in the sample.
@@ 4 Discussion
The corona properties we derived in Sect. 3.3 are roughly in agreement with what has been reported by other authors in the literature (Behar et al., 2015; Inoue & Doi, 2018), which raises two points. First, we note that the sizes inferred from the mm observations are consistently larger than the ones inferred from X-ray observations (Fabian et al., 2015) or microlensing (Rybak et al., 2025). Second, we note that the value we obtain of $\delta\approx 1.4$% is in agreement with the value required to explain the MeV background (Inoue et al., 2008, 2019). In addition, this value of $\delta$ corresponds to $\epsilon_{B}\sim 0.5$, leading to $\beta_{B}\sim 10$ (Sect. 2.1). We note that values of $\beta_{B}>1$ may challenge the ability of magnetic fields to sustain particle acceleration and heating in the corona (Inoue et al., 2024).
@@
Below, we provide an overview of the main limitations of our multi-component SED fitting, both from the modelling perspective and the observational aspects, in order to suggest possible ways to move forward.
@@
First, we highlight that we are using a one-zone coronal emission model. Even though the assumption of a homogeneous emitter is most likely an oversimplification, it still yields results that are currently consistent with the observations. It remains extremely difficult to constrain more complex models (e.g. Raginski & Laor, 2016) unless the optically thick part of the radio-mm SED is clearly detected. In most sources, this will require very deep and high-angular-resolution observations at frequencies below 100 GHz. Notably, the correlation between radio and X-ray luminosity is weaker at 22 GHz (Magno et al., 2025) than at 100 GHz (Ricci et al., 2023), consistent with the weaker coronal emission at 22 GHz predicted by our one-zone model. Another limitation of assuming a homogeneous spherical corona is that we cannot predict realistic images of it, although their characteristic sizes of below a few $\mu$as make them unresolvable with current facilities. As a final note on the coronae, we note that when observing very dust-obscured sources, we could expect absorption of its emission by dust at frequencies above $\nu_{\tau_{1}}$, although this depends on the dust being along our line of sight to the corona.
@@
Further limitations of the methodology used here include not accounting for an AME component from spinning dust grains, which could be relevant in sources that present a bump in the SED at $\sim$30 GHz (which is not the case for the sources in our sample). This component could be added at the expense of additional free parameters. Regarding the synchrotron component, we preferred to ignore the effects of f–f absorption (FFA) in this component due to the limited data at high angular resolution at low frequencies, although this possibility is available in our code (an application of a clumpy absorption medium was presented in Mutie et al., 2025). In this regard, it could be possible to adopt a more model-dependent approach in which, under certain assumptions about the diffuse gas distribution, the FFA is linked to the f–f emission, but this should be properly tested on a large sample of galaxies not necessarily limited to RQ AGN with corona emission. For this reason, we currently favour a more phenomenological approach in which the parameters that describe the FFA and f–f emission are not tied to further assumptions.
@@
Finally, we highlight the observational challenges involved in studying the corona emission of RQ AGN. From a fitting perspective, it is crucial to have a well-sampled SED from radio to FIR with consistent angular resolutions and $uv$-coverage. This ensures that the same scales are probed, although it is very challenging to obtain such a consistent dataset (so far only attempted in Mutie et al., 2025). High-resolution IR imaging is currently unavailable, which hinders our chances of constraining the dust component. The next generation VLA (ngVLA [^9] https://ngvla.nrao.edu) is very promising due to the extremely good angular resolution and sensitivity to frequencies up to $\sim$100 GHz that it will provide. In the mm to sub-mm domain, ALMA will remain the most suitable observatory.
One of the most difficult issues to address in the SED fitting is the corona variability on timescales of days. So far, the most extreme case of variability was found in IC 4329A, in which the mm flux density varied by a factor $\sim$3 and the inferred corona size by a factor of $\sim$2, while $\delta$ remained about constant (Shablovinskaya et al., 2024). Other sources have shown even less variability, and thus the corresponding change in the coronal parameters should be even smaller. However, this is difficult to assess due to the very limited studies of this phenomenon to date, and the lack of broad frequency coverage needed to get a good SED fit. In this regard, the prospects for the ALMA Wide Sensitivity Upgrade are particularly promising, as it will allow us to obtain a broad bandwidth in a single epoch, something particularly useful around $\sim$100 GHz (as shown in Shablovinskaya et al., 2024). Complementing this with simultaneous hard X-ray observations is ideal, as it can further reduce the uncertainty in the corona parameters $\tau_{\mathrm{T}}$ and $k\,T_{\mathrm{c}}$, although this should not be a deciding factor in most cases (Fig. 2).